OpenStudy (anonymous):

Celeste made a model of her grandfather’s fishing boat using a scale where 2 inches represents 3 feet. Her grandfather’s actual boat is 28 feet long. What is the length of Celeste’s model boat?

OpenStudy (anonymous):

the ratio looks like this 2/3=28/x, so 2x=84,then x= 42.

OpenStudy (anonymous):

the answer in my workbook is 12 in . Are you sure ?